# Physics-informed, boundary-constrained Gaussian process regression for the reconstruction of fluid flow fields

### v1

@@ 3.2 Constraints over parameterized sets
The following corollaries present two specific choices of ${\nu}$ that simplify Eq. (3.10), and are helpful for computing derivatives of this kernel in Sect. 4. In the first scenario, the geometry of $\Gamma$ has no other restriction than to be compact in ${\Omega}$. In the second scenario, we suppose $\Gamma$ to be a Lipschitz $(d-1)$-manifold in $\mathbb{R}^{d}$.
@@
In the context of Proposition 3.1, consider $\Gamma$ as a compact manifold contained in ${\Omega}$ such that we can define a parameterization ${\boldsymbol{\gamma}}$ with domain ${\mathcal{P}}\subset\mathbb{R}^{p}$ (for an integer $p$) that is $({\mathcal{P}},\Gamma)$-measurable and such that $\Gamma=\{{\boldsymbol{\gamma}}({\boldsymbol{s}})=(\gamma_{1}({\boldsymbol{s}}),\dots,\gamma_{d}({\boldsymbol{s}}))^{\mathsf{T}},\ {\boldsymbol{s}}\in{\mathcal{P}}\}$. Then, by choosing an adequate measure ${\nu}$ for the Karhunen-Loève expansion of ${Z}$ over $\Gamma$, we have:
@@
In the context of Proposition 3.1, consider $\Gamma$ as a compact Lipschitz $(d-1)$-manifold contained in ${\Omega}$ such that we can define an injective Lipschitz parameterization ${\boldsymbol{\gamma}}$ with domain ${\mathcal{P}}\subset\mathbb{R}^{p}$ (for an integer $p$) such that $\Gamma=\{{\boldsymbol{\gamma}}({\boldsymbol{s}})=(\gamma_{1}({\boldsymbol{s}}),\ldots,\gamma_{d}({\boldsymbol{s}}))^{\mathsf{T}},\ {\boldsymbol{s}}\in{\mathcal{P}}\}$. Then, by choosing an adequate measure ${\nu}$ for the Karhunen-Loève expansion of ${Z}$ over $\Gamma$, we have:
@@
We actually choose ${\nu}$ as the surface measure ${\mathcal{H}}^{d-1}$ on $\mathbb{R}^{d}$ normalized by ${\mathcal{H}}^{d-1}(\Gamma)$, the $(p-1)$-dimensional measure of the manifold. The Lipschitz assumption on the curve $\Gamma$ implies that the Jacobian of ${\boldsymbol{\gamma}}$ is defined $a.e.$ on ${\mathcal{P}}$. Then, by the change of variables formula on submanifolds from $\mathbb{R}^{p}$ to $\mathbb{R}^{d}$ (see Section 3.3 in [15]), the integral in Eq. (3.3) reads:
@@ 3.3 A numerical method for approximating constrained kernels
In this section, we describe a numerical method to compute kernels for BCGPs, as defined by Eq. (3.10) in the case where we consider a curve $\Gamma\subset\mathbb{R}^{d}$ with a Lipschitz parameterization ${\boldsymbol{\gamma}}:{\mathcal{P}}\rightarrow\Gamma$ and ${\mathcal{P}}\subset\mathbb{R}^{p}$. For simplicity, we consider $p=1$ and ${\mathcal{P}}=[0,2\pi]$. As before, we denote by ${{Z}_{0}}$ the associated BCGP, which is computed from an arbitrary base GP ${Z}$ over the bounded domain ${\Omega}$, with continuous kernel ${G}$. Since this continuous setting is described in terms of a spectral expansion of the base kernel, we perform a truncation of this spectrum. However, it is worth to remark that this numerical procedure does not represent a discrete enforcement of the boundary constraint (3.6) at collocation points, but it is instead an approximation of the geometry of the compact set verifying (3.6) at all points of the geometry.
We outline how the boundary-constrained kernel ${{G}_{0}}$ obtained by Eq. (3.10) is computed through a carefully crafted approximation in terms of mode truncation and numerical integration. In order to completely define ${{G}_{0}}$, we seek to approximate the spectral pairs $\smash{(\lambda_{n},{\phi}_{n})}$ by discrete counterparts. By definition, these pairs are obtained from the integral representation with respect to the base kernel ${G}$:
@@
where ${\boldsymbol{G}}$ is the symmetric matrix with elements $[{\boldsymbol{G}}]_{ij}={G}\left({\boldsymbol{\gamma}}(s_{i}),{\boldsymbol{\gamma}}(s_{j})\right)$, $i,j=0,\ldots,I-1$; ${\boldsymbol{\Lambda}}=\smash{\operatorname{diag}(\lambda^{(I)}_{0},\lambda^{(I)}_{1},\ldots,\lambda^{(I)}_{I-1})}\in\smash{\mathcal{M}_{I}(\mathbb{R}^{+})}$ is the diagonal matrix of eigenvalues; and ${\boldsymbol{E}}=\smash{(\boldsymbol{{\phi}}^{(I)}_{0},\ldots\boldsymbol{{\phi}}^{(I)}_{I-1})}\in\mathcal{M}_{I}(\mathbb{R})$ is the matrix of orthogonal eigenvectors. By using the square root of ${\boldsymbol{H}}$, we can transform Eq. (3.17) to the classical eigenvalue problem:
@@
In practice, given an arbitrary continuous base kernel ${G}$, we must consider a spectral truncation of ${{G}_{0}}$. In order to quantify the convergence of the spectral series, we define the spectral accuracy over $\Gamma$ as:
@@
Input: Base kernel ${G}$ with fixed hyperparameters, number of integration intervals $I$, profile boundary nodes ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{k}_{\Gamma}\}_{k=0}^{I-1}}$, measure weights ${\boldsymbol{H}}$, spectral accuracy bound $\delta$
@@
7Compute the minimal number of spectral modes $J+1$ satisfying the condition (3.20) given $I$ and $\delta$;
@@ 3.4 Benefits and limitations with respect to discrete constraints
On the other hand, the inversion cost with the proposed method is of order $\text{O}\smash{(N^{3})}$ since no observation points are required in the compact set $\Gamma$. Even though the evaluation of the BCGP kernel ${{G}_{0}}$ is more expensive in comparison to that of ${G}$ (see Table 1), the computational complexity of this procedure is $\text{O}\smash{(N(J+1)I)}$, for $J+1$ being the number of spectral modes and $I+1$ being the number of selected nodes in ${\mathcal{P}}$ for the approximation of the eigenvalue problem. Accordingly, the total computational cost of using BCGPs is not larger than $\text{O}\smash{(N^{3})}$ since in practice $J\ll N$. Furthermore, the main advantage of using BCGPs relies on the fact that the constraint (3.6) is satisfied continuously, where the approximation procedure is not a spatial discretization but a discretization of the spectrum defining the geometry of $\Gamma$.
@@
Table 1: Comparison of the complexity of two approaches for the enforcement of the Dirichlet boundary condition (3.6) on the posterior distribution from the base GP ${Z}$. As a reference, we consider the complexity of a single evaluation of kernel ${G}$ of order O$(1)$.
@@ 4.1 Fluid dynamics context
The external boundary $\partial{\Omega}$ of ${\Omega}$ excluding the boundary of the profile $\Gamma$ is the rectangle from 00 to $L_{1}$ in the horizontal (streamwise) direction and from 00 to $L_{2}$ in the vertical one. The boundary conditions over $\partial{\Omega}=\partial{\Omega}_{\text{in}}\cup\partial{\Omega}_{\text{out}}\cup\partial{\Omega}_{\text{w}}$ are:
@@
    ${\boldsymbol{u}}({\boldsymbol{x}},t)=\begin{cases}{{\boldsymbol{u}}_{\text{in}}}({\boldsymbol{x}},t)\ ,&\text{for }{\boldsymbol{x}}\in\partial{\Omega}_{\text{in}}=\left\{0\right\}\times(0,L_{2})\,,\\ {{\boldsymbol{u}}_{\text{out}}}({\boldsymbol{x}},t)\ ,&\text{for }{\boldsymbol{x}}\in\partial{\Omega}_{\text{out}}=\left\{L_{1}\right\}\times(0,L_{2})\,.\\ \end{cases}$ (4.3)
@@ 4.2 Physics-informed, boundary-constrained GP
More precisely, we consider an arbitrary GP ${Z}$ indexed on the spatial domain ${\Omega}$ with a covariance kernel ${G}$ belonging to $\mathcal{C}^{3}({\Omega}\times{\Omega})$. Since the profile boundary $\Gamma$ is a compact set inside ${\Omega}$, we model ${\psi}$ by a BCGP defined from this base GP ${Z}$. That is, we consider ${\psi}={{Z}_{0}}\sim{\mathcal{GP}}(0,{{G}_{0}})$ as the BCGP associated to ${Z}$ through the procedure of Proposition (3.1), for ${{G}_{0}}$ being the constrained kernel defined in Eq. (3.10). It follows that the flow velocity can be modelled as the zero-mean vector-valued GP:
@@
for ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$, with ${\textbf{curl}}_{\boldsymbol{x}}=\smash{{\boldsymbol{\nabla}}_{\boldsymbol{x}}^{\perp}}=\smash{(-\partial_{x_{2}},\partial_{x_{1}})^{\mathsf{T}}}$ and ${\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}=\smash{{\boldsymbol{\nabla}}_{{\boldsymbol{x}}^{\prime}}^{\perp}}=\smash{(-\partial_{x^{\prime}_{2}},\partial_{x^{\prime}_{1}})^{\mathsf{T}}}$ acting on the first and the second entries of ${{G}_{0}}$, respectively. This kernel is well-defined pointwise thanks to the regularity of ${G}$ [18, 33]. Then we observe that in this case the GP ${\boldsymbol{u}}={\textbf{curl}}\,{{Z}_{0}}$ satisfies for each ${\boldsymbol{x}}\in{\Omega}$:
@@ 4.3 GPR observation design
0 $L_{1}$ $L_{2}$ profile:${\boldsymbol{u}}\cdot{\boldsymbol{n}}=0$top wall: ${{\boldsymbol{u}}_{\text{w}}}$bottom wall: ${{\boldsymbol{u}}_{\text{w}}}$${{\boldsymbol{u}}_{\text{in}}}$${{\boldsymbol{u}}_{\text{out}}}$
@@
$\begin{split}{{\omega}^{\star}}({\boldsymbol{x}},t)&=\Delta_{\boldsymbol{x}}({\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}\,{{G}_{0}}({\boldsymbol{x}},{\boldsymbol{X}})\cdot{\boldsymbol{\beta}_{0}}(t))\\ &={\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}\Delta_{\boldsymbol{x}}{{G}_{0}}({\boldsymbol{x}},{\boldsymbol{X}})\cdot{\boldsymbol{\beta}_{0}}(t)\,,\quad{\boldsymbol{x}}\in{\Omega}\,,\end{split}$ (4.13)
@@ 4.4 Derivatives of a constrained kernel
We see from Eqs. (4.9), (4.11) and (4.13) that derivatives of the BCGP kernels are needed in the proposed approach. Here we describe how they may be computed. Let us first notice that in the case the kernel ${G}$ is regular enough, the uniform convergence of the series in Eq. (3.10) allows us to compute the partial derivatives of ${{G}_{0}}$ up to the same regularity as that of ${G}$. Then for multi-indices $\smash{\alpha,\beta\in\mathbb{N}^{2}}$ and ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$, such that $\smash{\partial_{\boldsymbol{x}}^{\alpha}}=\smash{\partial_{x_{1}}^{\alpha_{1}}\partial_{x_{2}}^{\alpha_{2}}}$ and similarly for $\smash{\partial^{\beta}_{{\boldsymbol{x}}^{\prime}}}$, the partial derivatives of ${{G}_{0}}$ are computed directly from Eq. (3.10) as:
@@
Consider the case of a parameterized curve $\Gamma$ of class $C^{k}$, denoting as ${\boldsymbol{\gamma}}$ its regular diffeomorphism. Then, as a consequence of the Mercer decomposition of ${G}$ along $\Gamma$, we have that:
@@
and for any derivatives in the curve direction with multi-indices $\alpha$ and $\beta$ of order $k$. Nevertheless, it is worth to notice that given multi-indices $\alpha,\beta$, the derivatives $\smash{\partial^{\alpha}_{\boldsymbol{x}}\otimes\partial^{\beta}_{{\boldsymbol{x}}^{\prime}}\,{{G}_{0}}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})}$ are not necessarily null for all ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$ since the spectral expansion of ${G}$ through the pairs $\smash{(\lambda_{n},{\phi}_{n})_{n=0}^{+\infty}}$ is obtained specifically from its restriction to $\Gamma$, and not over all ${\Omega}$.
@@
Input: Base kernel ${G}$ with fixed hyperparameters, evaluation points ${{\boldsymbol{X}}}=\smash{\{{\boldsymbol{x}}_{i}\}_{i=1}^{N_{{\boldsymbol{X}}}}}$ and ${{\boldsymbol{X}}^{\prime}}=\smash{\{{\boldsymbol{x}}^{\prime}_{j}\}_{j=1}^{N_{{\boldsymbol{X}}^{\prime}}}}$, spectral factor matrix ${\boldsymbol{S}}$, profile boundary nodes ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{k}_{\Gamma}\}_{k=0}^{I-1}}$, derivative order $M$

### v2

@@ 3.2 Constraints over parameterized sets
The following corollaries present two specific choices of $\nu$} that simplify Eq. (3.10), and are helpful for computing derivatives of this kernel in Sect. 4. In the first scenario, the geometry of $\Gamma$ has no other restriction than to be compact in $\Omega$}. In the second scenario, we suppose $\Gamma$ to be a Lipschitz (d1)(d-1)-manifold in $\mathbb{R}^{d}$.
@@
In the context of Proposition 3.1, consider $\Gamma$ as a compact manifold contained in $\Omega$} such that we can define a parameterization $\boldsymbol{\gamma}$} with domain ${\mathcal{P}}\subset\mathbb{R}^{p}$ (for an integer $p$) that is $({\mathcal{P}},\Gamma)$-measurable and such that $\Gamma=\{{\boldsymbol{\gamma}}({\boldsymbol{s}})=(\gamma_{1}({\boldsymbol{s}}),\dots,\gamma_{d}({\boldsymbol{s}}))^{\mathsf{T}},\ {\boldsymbol{s}}\in{\mathcal{P}}\}$. Then, by choosing an adequate measure $\nu$} for the Karhunen-Loève expansion of Z{Z} over $\Gamma$, we have:
@@
In the context of Proposition 3.1, consider $\Gamma$ as a compact Lipschitz (d1)(d-1)-manifold contained in $\Omega$} such that we can define an injective Lipschitz parameterization $\boldsymbol{\gamma}$} with domain ${\mathcal{P}}\subset\mathbb{R}^{p}$ (for an integer $p$) such that $\Gamma=\{{\boldsymbol{\gamma}}({\boldsymbol{s}})=(\gamma_{1}({\boldsymbol{s}}),\ldots,\gamma_{d}({\boldsymbol{s}}))^{\mathsf{T}},\ {\boldsymbol{s}}\in{\mathcal{P}}\}$. Then, by choosing an adequate measure $\nu$} for the Karhunen-Loève expansion of Z{Z} over $\Gamma$, we have:
@@
We actually choose $\nu$} as the surface measure ${\mathcal{H}}^{d-1}$ on $\mathbb{R}^{d}$ normalized by ${\mathcal{H}}^{d-1}(\Gamma)$, the (p1)(p-1)-dimensional measure of the manifold. The Lipschitz assumption on the curve $\Gamma$ implies that the Jacobian of $\boldsymbol{\gamma}$} is defined $a.e.$ on $\mathcal{P}$}. Then, by the change of variables formula on submanifolds from $\mathbb{R}^{p}$ to $\mathbb{R}^{d}$ (see Section 3.3 in [15]), the integral in Eq. (3.3) reads:
@@ 3.3 A numerical method for approximating constrained kernels
In this section, we describe a numerical method to compute kernels for BCGPs, as defined by Eq. (3.10) in the case where we consider a curve $\Gamma\subset\mathbb{R}^{d}$ with a Lipschitz parameterization ${\boldsymbol{\gamma}}:{\mathcal{P}}\rightarrow\Gamma$ and ${\mathcal{P}}\subset\mathbb{R}^{p}$. For simplicity, we consider $p=1$ and ${\mathcal{P}}=[0,2\pi]$. As before, we denote by ${Z}_{0}$} the associated BCGP, which is computed from an arbitrary base GP Z{Z} over the bounded domain $\Omega$}, with continuous kernel G{G}. Since this continuous setting is described in terms of a spectral expansion of the base kernel, we perform a truncation of this spectrum. However, it is worth to remark that this numerical procedure does not represent a discrete enforcement of the boundary constraint (3.6) at collocation points, but it is instead an approximation of the geometry of the compact set verifying (3.6) at all points of the geometry.
We outline how the boundary-constrained kernel ${G}_{0}$} obtained by Eq. (3.10) is computed through a carefully crafted approximation in terms of mode truncation and numerical integration. In order to completely define ${G}_{0}$}, we seek to approximate the spectral pairs $\smash{(\lambda_{n},{\phi}_{n})}$ by discrete counterparts. By definition, these pairs are obtained from the integral representation with respect to the base kernel G{G}:
@@
where $\boldsymbol{G}$} is the symmetric matrix with elements ${\boldsymbol{G}}]_{ij}={G}\left({\boldsymbol{\gamma}}(s_{i}),{\boldsymbol{\gamma}}(s_{j})\right)$, $i,j=0,\ldots,I-1$; ${\boldsymbol{\Lambda}}=\smash{\operatorname{diag}(\lambda^{(I)}_{0},\lambda^{(I)}_{1},\ldots,\lambda^{(I)}_{I-1})}\in\smash{\mathcal{M}_{I}(\mathbb{R}^{+})}$ is the diagonal matrix of eigenvalues; and ${\boldsymbol{E}}=\smash{(\boldsymbol{{\phi}}^{(I)}_{0},\ldots\boldsymbol{{\phi}}^{(I)}_{I-1})}\in\mathcal{M}_{I}(\mathbb{R})$ is the matrix of orthogonal eigenvectors. By using the square root of $\boldsymbol{H}$}, we can transform Eq. (3.17) to the classical eigenvalue problem:
@@
In practice, given an arbitrary continuous base kernel G{G}, we must consider a spectral truncation of ${G}_{0}$}. In order to quantify the convergence of the spectral series, we define the spectral accuracy over $\Gamma$ as:
@@
Input: Base kernel G{G} with fixed hyperparameters, number of integration intervals II, profile boundary nodes ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{k}_{\Gamma}\}_{k=0}^{I-1}}$, measure weights $\boldsymbol{H}$}, spectral accuracy bound $\delta$
@@
7Compute the minimal number of spectral modes $J+1$ satisfying the condition (3.20) given II and $\delta$;
@@ 3.4 Benefits and limitations with respect to discrete constraints
On the other hand, the inversion cost with the proposed method is of order $\text{O}\smash{(N^{3})}$ since no observation points are required in the compact set $\Gamma$. Even though the evaluation of the BCGP kernel ${G}_{0}$} is more expensive in comparison to that of G{G} (see Table 1), the computational complexity of this procedure is $\text{O}\smash{(N(J+1)I)}$, for $J+1$ being the number of spectral modes and $I+1$ being the number of selected nodes in $\mathcal{P}$} for the approximation of the eigenvalue problem. Accordingly, the total computational cost of using BCGPs is not larger than $\text{O}\smash{(N^{3})}$ since in practice $J\ll N$. Furthermore, the main advantage of using BCGPs relies on the fact that the constraint (3.6) is satisfied continuously, where the approximation procedure is not a spatial discretization but a discretization of the spectrum defining the geometry of $\Gamma$.
@@
Table 1: Comparison of the complexity of two approaches for the enforcement of the Dirichlet boundary condition (3.6) on the posterior distribution from the base GP Z{Z}. As a reference, we consider the complexity of a single evaluation of kernel G{G} of order O(1)(1).
@@ 4.1 Fluid dynamics context
The external boundary $\partial{\Omega}$ of $\Omega$} excluding the boundary of the profile $\Gamma$ is the rectangle from 0 to $L_{1}$ in the horizontal (streamwise) direction and from 0 to $L_{2}$ in the vertical one. The boundary conditions over $\partial{\Omega}=\partial{\Omega}_{\text{in}}\cup\partial{\Omega}_{\text{out}}\cup\partial{\Omega}_{\text{w}}$ are:
@@
[…]
@@ 4.2 Physics-informed, boundary-constrained GP
More precisely, we consider an arbitrary GP Z{Z} indexed on the spatial domain $\Omega$} with a covariance kernel G{G} belonging to $\mathcal{C}^{3}({\Omega}\times{\Omega})$. Since the profile boundary $\Gamma$ is a compact set inside $\Omega$}, we model $\psi$} by a BCGP defined from this base GP Z{Z}. That is, we consider ${\psi}={{Z}_{0}}\sim{\mathcal{GP}}(0,{{G}_{0}})$ as the BCGP associated to Z{Z} through the procedure of Proposition (3.1), for ${G}_{0}$} being the constrained kernel defined in Eq. (3.10). It follows that the flow velocity can be modelled as the zero-mean vector-valued GP:
@@
for ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$, with ${\textbf{curl}}_{\boldsymbol{x}}=\smash{{\boldsymbol{\nabla}}_{\boldsymbol{x}}^{\perp}}=\smash{(-\partial_{x_{2}},\partial_{x_{1}})^{\mathsf{T}}}$ and ${\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}=\smash{{\boldsymbol{\nabla}}_{{\boldsymbol{x}}^{\prime}}^{\perp}}=\smash{(-\partial_{x^{\prime}_{2}},\partial_{x^{\prime}_{1}})^{\mathsf{T}}}$ acting on the first and the second entries of ${G}_{0}$}, respectively. This kernel is well-defined pointwise thanks to the regularity of G{G} [18, 34]. Then we observe that in this case the GP ${\boldsymbol{u}}={\textbf{curl}}\,{{Z}_{0}}$ satisfies for each ${\boldsymbol{x}}\in{\Omega}$:
@@ 4.3 GPR observation design
[…]
@@
[…]
@@ 4.4 Derivatives of a constrained kernel
We see from Eqs. (4.9), (4.11) and (4.13) that derivatives of the BCGP kernels are needed in the proposed approach. Here we describe how they may be computed. Let us first notice that in the case the kernel G{G} is regular enough, the uniform convergence of the series in Eq. (3.10) allows us to compute the partial derivatives of ${G}_{0}$} up to the same regularity as that of G{G}. Then for multi-indices $\smash{\alpha,\beta\in\mathbb{N}^{2}}$ and ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$, such that $\smash{\partial_{\boldsymbol{x}}^{\alpha}}=\smash{\partial_{x_{1}}^{\alpha_{1}}\partial_{x_{2}}^{\alpha_{2}}}$ and similarly for $\smash{\partial^{\beta}_{{\boldsymbol{x}}^{\prime}}}$, the partial derivatives of ${G}_{0}$} are computed directly from Eq. (3.10) as:
@@
Consider the case of a parameterized curve $\Gamma$ of class $C^{k}$, denoting as $\boldsymbol{\gamma}$} its regular diffeomorphism. Then, as a consequence of the Mercer decomposition of G{G} along $\Gamma$, we have that:
@@
and for any derivatives in the curve direction with multi-indices $\alpha$ and $\beta$ of order $k$. Nevertheless, it is worth to notice that given multi-indices $\alpha,\beta$, the derivatives $\smash{\partial^{\alpha}_{\boldsymbol{x}}\otimes\partial^{\beta}_{{\boldsymbol{x}}^{\prime}}\,{{G}_{0}}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})}$ are not necessarily null for all ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}$ since the spectral expansion of G{G} through the pairs $\smash{(\lambda_{n},{\phi}_{n})_{n=0}^{+\infty}}$ is obtained specifically from its restriction to $\Gamma$, and not over all $\Omega$}.
@@
Input: Base kernel G{G} with fixed hyperparameters, evaluation points ${{\boldsymbol{X}}}=\smash{\{{\boldsymbol{x}}_{i}\}_{i=1}^{N_{{\boldsymbol{X}}}}}$ and ${{\boldsymbol{X}}^{\prime}}=\smash{\{{\boldsymbol{x}}^{\prime}_{j}\}_{j=1}^{N_{{\boldsymbol{X}}^{\prime}}}}$, spectral factor matrix $\boldsymbol{S}$}, profile boundary nodes ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{k}_{\Gamma}\}_{k=0}^{I-1}}$, derivative order $M$